OpenStudy (anonymous):

What is the 41st term of the arithmetic sequence where a1 = 16 and a15 = –40 ?

OpenStudy (anonymous):

-144

OpenStudy (cwrw238):

common difference =( a15 - a1 ) / 14 = -56 /14 = -4 41st term = 16 + 40(-4)

OpenStudy (cwrw238):

swagdaddy - please note the method - we like to teach you how not just give the answer

OpenStudy (anonymous):

i like to give answers

OpenStudy (cwrw238):

well you ought to read the open study rules.....